OpenStudy (anonymous):

whats is the exact value of csc^-1(2)

OpenStudy (tkhunny):

So, \(A = \csc^{-1}(2)\rightarrow \csc(A) = 2 \rightarrow \dfrac{1}{\sin(A)} = 2 \rightarrow \sin(A) = 1/2\) That looks easier.